OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). who knows the answer to this ?

OpenStudy (amistre64):

well, id hope you do .....

OpenStudy (amistre64):

f(x) = -4x + 7 let x=g f(g) = -4g + 7 but then g = 10x - 6 f(g(x)) = -4(10x - 6) + 7